Using ∈-δ arguments, prove that


Lim (x^3+1) = 9/8


x→1/2

1
Expert's answer
2021-11-23T13:25:50-0500

Given "\epsilon" > 0, we have to find "\delta>0" such that

if "0< |x-\frac{1}{2}|< \delta" then | x³+1"-\frac{9}{8}|" |<"\epsilon"

Now | x³+1 "-\frac{9}{8}|| = |x³-\frac{1}{8}|"

= |(x "-\frac{1}{2})(x²+\frac{1}{2}x+\frac{1}{4})|"

If |x"-\frac{1}{2}|<\frac{1}{2}" that means "-\frac{1}{2} <x-\frac{1}{2} < \frac{1}{2}" then

0 < x < 1 <=> "x²+\frac{1}{2}x +\frac{1}{4} < 1²+ \frac{1}{2} +\frac{1}{4}= \frac{7}{4}"

Therefore |x³+1"-\frac{9}{8}| = |x- \frac{1}{2}|" (x²+"-\frac{1}{2})(x²+\frac{1}{2}x+\frac{1}{4})< \frac{7}{4}|x-\frac{1}{2}|"

Let us choose "\delta =" min { "\frac{1}{2} , \frac{4\epsilon}{7}" }

then "0< |x-\frac{1}{2}|< \delta" => |x³+1"-\frac{9}{8}|| < \frac{4\epsilon}{7}.\frac{7}{4}=\epsilon"

Thus by the "\delta-\epsilon" method

"lim_{x->\frac{1}{2}}(x³+1) = \frac{9}{8}"